### v1

# IntactKV: Improving Large Language Model Quantization by Keeping Pivot Tokens Intact

Abstract
Large language models (LLMs) excel in natural language processing but demand intensive computation. To mitigate this, various quantization methods have been explored, yet they compromise LLM performance. This paper unveils a previously overlooked type of outlier in LLMs. Such outliers are found to allocate most of the attention scores on initial tokens of input, termed as pivot tokens, which is crucial to the performance of quantized LLMs. Given that, we propose IntactKV to generate the KV cache of pivot tokens losslessly from the full-precision model. The approach is simple and easy to combine with existing quantization solutions. Besides, IntactKV can be calibrated as additional LLM parameters to boost the quantized LLMs further. Mathematical analysis also proves that IntactKV effectively reduces the upper bound of quantization error. Empirical results show that IntactKV brings consistent improvement and achieves lossless weight-only INT4 quantization on various downstream tasks, leading to the new state-of-the-art for LLM quantization.
IntactKV: Improving Large Language Model Quantization by
Keeping Pivot Tokens Intact
[…]
1 Introduction
Large language models (LLMs) have achieved remarkable progress in various tasks and benchmarks in natural language processing Brown et al. (2020); Bubeck et al. (2023); Touvron et al. (2023a); Team et al. (2023). Nonetheless, the rise of LLMs also increases computational intensity and memory requirements. This motivates various research to decrease the inference cost of LLMs, e.g., quantizaiton Frantar et al. (2022); Shao et al. (2024); Lin et al. (2023), pruning Frantar and Alistarh (2023); Liu et al. (2023b); Sun et al. (2023), and speculative decoding Chen et al. (2023); Leviathan et al. (2023); Cai et al. (2024), e.t.c.
Among these methods, network quantization converts the network parameters or activations from floating-point to fixed-point formats, which is a popular technique to reduce the model size and computational resources. Nevertheless, quantization inevitably affects the performance of LLMs. The leading cause comes from the outliers in LLM activations, which are sensitive to network quantization Dettmers et al. (2022); Xiao et al. (2023); Lin et al. (2023). As workarounds, there are efforts to either use mixed-precision formats Dettmers et al. (2022) or re-scale network weights of the outlier channels Lin et al. (2023). These methods are all built based on the premise that outliers persist in fixed channels across all tokens. However, we find this is not the case for all outliers in LLMs.
In this paper, we discover a new type of outlier that is overlooked by previous quantization methods. These outliers exhibit extremely high values at only the [BOS] and some other common tokens (e.g., “,” and “.”) at the beginning of the input, which is referred to as pivot tokens. We find the extreme values of these outliers make the self-attention concentrate on the pivot tokens, leaving the rest of the tokens untouched. This is also known as attention sinks Xiao et al. (2024), which is critical to the model performance Xiao et al. (2024); Bondarenko et al. (2023). The effect of quantization on these pivot tokens should be carefully studied to improve the quantized LLMs.
Figure 1: Visualizations of Transformer output and attention scores of LLaMA-30B and LLaMA-2-7B. Observations: (1) There are token-specific outliers that can be orders of magnitudes larger than the rest of the tokens (enlarged in the box). Such tokens occur at the [BOS] token, the 28th token "" in LLaMA-30B and 13th token "." in LLaMA-2-7B, which are referred to as pivot tokens; (2) These outliers over pivot tokens make the attention scores concentrated on themselves, which are likely to be affected by quantization. More details can be found in Appendix C.1.
Figure 2: The mean squared error (MSE) of the last Transformer layer and attention layers w.r.t. the varying sizes of IntactKV. Observations: (1) The MSE continues to drop as the size of IntactKV increases. (2) Including the pivot tokens’ KV cache in IntactKV leads to the most significant decrease in the quantization loss, demonstrating the importance of the pivot tokens’ KV cache. More experiment details can be found in Appendix D.
Towards that end, we are motivated to propose IntactKV, a simple strategy that is orthogonal to most existing quantization solutions. The key idea behind IntactKV is to generate the KV cache of pivot tokens from the full-precision model. By keeping the KV cache of pivot tokens intact, quantization error accumulated on the output of self-attention will be effectively alleviated in the rest of the decoding steps. Moreover, IntactKV can also serve as extra trainable parameters in addition to the LLM backbone. The calibration process of IntactKV follows the convention of PTQ Bai et al. (2022); Frantar et al. (2022); Lin et al. (2023), which further decreases the quantization error. To get more insights from IntactKV, we also provide mathematical analysis and the results show that IntactKV can effectively lower the upper bound of quantization error.
Empirical results show that IntactKV consistently improves the capability of quantized models on various open-sourced LLMs (e.g., LLaMA and Vicuna) and different downstream tasks such as C4, MMLU, commonsense QA, and MT-bench. When armed with AWQ Lin et al. (2023), IntactKV achieves new state-of-the-art quantization results, e.g., lossless INT4 weight-only quantization for Vicuna-v1.5 on commonsense QA tasks. Moreover, fine-tuning IntactKV with INT4 quantization even matches the full-precision model on aligning with human preference, as evaluated by GPT-4 Bubeck et al. (2023) on MT-bench.
2 Motivation
2.1 Preliminaries on LLM Quantization
Network quantization is popularly studied in the literature of efficient LLMs Frantar et al. (2022); Lin et al. (2023); Shao et al. (2024). It allows larger throughput by reducing the model size and leads to practical inference speedup. Given the full-precision weight $\mathbf{w}$, quantization aims to convert it to the low-bit representation $\hat{\mathbf{w}}$. The general $b$-bit uniform quantization $\mathcal{Q}_{b}(\cdot)$ can be represented as
$\hat{\mathbf{w}}=\mathcal{Q}_{b}(\mathbf{w})=s\cdot\Pi_{\Omega(b)}(\mathbf{w}/s),$ (1)
where $s$ is the quantization step size, and $\Pi_{\Omega(b)}$ is the projection function onto the set of $b$-bit integers $\Omega(b)=\{0,1,...,2^{b}-1\}$. While we mainly focus on weight-only quantization, Equation 1 can be similarly used to quantize activations and KV cache of LLMs to increase the inference throughput Xiao et al. (2023); Shao et al. (2024); Hooper et al. (2024).
Following most existing works in LLM quantization, we focus on post-training quantization (PTQ) Frantar et al. (2022); Lin et al. (2023), since it does not introduce extra training overhead as those in quantization-aware training (QAT) Liu et al. (2023a); Li et al. (2024). Quantization inevitably downgrades LLMs in low-bit settings, where the outliers in quantized LLMs are found to be the cause of the deterioration Dettmers et al. (2022). In the next, we study the details of how these outliers affect the LLM quantization.
Figure 3: The overview of the proposed IntactKV applied for the supervised fine-tuned LLM. The full-precision model takes the system prompt as input and generates the IntactKV losslessly as the prefix concatenated with the rest of the KV cache of quantized LLMs. IntactKV can be further calibrated by minimizing the mean square error $\mathcal{L}$ between the full-precision and quantized LLM.
2.2 Revisiting Outliers in LLMs
We discover a new type of outlier that is specific to particular tokens, which leads the attention sink Xiao et al. (2024) that is critical to the performance of LLMs.
A New Variant of Outlier.
Different from the outliers that persist in several fixed channels across different tokens (Dettmers et al., 2022; Xiao et al., 2023; Lin et al., 2023), we find a new variant of outlier that is specific to some initial tokens of the input sequence. By visualizing the activation of Transformer layer output in Figure 0(a) and Figure 0(b), there exist peaks with magnitudes over 1e3. These outliers can be hundreds of times larger than the previous outliers that persist in fixed channels across all tokens, as enlarged in Figure 0(a) and Figure 0(b). More visualizations can be found in Appendix C. It is found that such huge outliers usually occur at the [BOS] token and some other uninformative initial tokens (e.g., "." or ",") at particular channels, regardless of the rest of the input sequence. We thus name these tokens pivot tokens given their dominating values in the activation. Recently, a concurrent work (Sun et al., 2024) also discovers such outliers with more detailed studies.
Pivot Tokens Exhibit Attention Sinks.
We hypothesize that the outliers over these pivot tokens may propagate to queries and keys in the self-attention. Consequently, the attention scores will be concentrated on these pivot tokens than the rest ones, a.k.a attention sinks Xiao et al. (2024). To verify the hypothesis, we plot the attention scores in Figure 0(c) and Figure 0(d). It can be found that the pivot tokens indeed dominate the attention scores, especially for the first token (i.e., [BOS] ). This corresponds to the observations in attention sinks (Xiao et al., 2024), which are empirically verified to be critical to the model performance. The recent study by (Bondarenko et al., 2023) also shows that concentrating on these tokens naturally helps the attention head do nothing but simply a partial update of the residual. In the decoding stage of LLMs, all generated tokens need to interact with pivot tokens through self-attention. However, as mentioned in Section 2.1, network quantization would inevitably distort the output from the full-precision model. The concentrated scores of pivot tokens thus can be further deviated by quantization, which downgrades the model performance.
3 Method
In this section, we introduce IntactKV, a simple and easy-to-implement method to improve the quantized LLMs. The key idea behind this is to keep the KV cache of the pivot tokens intact, i.e., without any distortion raised by quantization. An overview of our method can be found in Figure 3.
3.1 Preserving the KV Cache of Pivot Tokens
According to Section 2.2, the attention sinks of pivot tokens are likely to deteriorate by quantization. To alleviate this issue, we propose IntactKV, a simple and effective strategy to keep these pivot tokens intact. Specifically, as illustrated in Figure 2(a), we leverage the full-precision LLM to generate the lossless KV cache of pivot tokens, which is saved offline. The quantized LLM then loads IntactKV as the prefix to concatenate with the rest of the KV cache and continues with the regular auto-regressive decoding. The pseudo code of the inference scheme with IntactKV is presented in Figure 2(b).
In order to study the benefits of IntactKV, we conduct a preliminary test on the mean square error (MSE) of the attention and transformer layer output. From Figure 2, it is natural that the increasing size of IntactKV gives the monotonically decreasing MSE on both the attention and transformer layers. More importantly, it is found the same tokens in Section 2.2 (e.g., [BOS] and other delimiter tokens) give the most significant decrease on the MSE, which demonstrates the importance of their KV cache. This aligns with the observations in Figure 1 that pivot tokens exhibit outliers with extreme values and attention sinks.
The Choice of Pivot Tokens and IntactKV.
It is the key design to choose the pivot tokens and the associated IntactKV. Given the observations in Figure 2, one can naively pick pivot tokens with the most MSE reduction for IntactKV. However, this is in fact not the case. Since IntactKV acts as the prefix to the KV cache of quantized LLMs, it must start from the very first token, and be consecutive in length. This ensures it to be input agnostic, and the full-precision LLMs can be safely discarded once IntactKV is generated. Next, we provide practical solutions to this problem for different LLMs.
  • For pre-trained LLMs, we propose the IntactKV of size one that only contains [BOS] KV cache. It is a convention to prepend [BOS] to the input of pre-trained LLMs. Moreover, as illustrated in Section 2, [BOS] is the pivot token with extreme outlier and attention scores. Besides, the KV cache of [BOS] has a great impact on the MSE of the quantized model. Employing a lossless [BOS] KV cache is thus believed to decrease the quantization loss.
  • For supervised fine-tuned (SFT) models, when the input follows the system prompt, we argue that extending IntactKV to the same length of the system prompt can further improve quantized LLMs. In addition to [BOS], other tokens appearing at the beginning of the input sequence also have the potential to serve as pivot tokens (see Figure 1). The system prompt is usually prepended to the input, which allows it to cover more pivot tokens. As shown in Figure 2, remedying the quantization error of these pivot tokens’ KV cache can be helpful to compensate for the quantization error. We find that for Vicuna models, system prompt is enough to cover all the pivot tokens, more details can be found in Appendix C.3.
Overhead of IntactKV.
Finally, we highlight that IntactKV does not introduce extra latency overhead during inference. Besides, as IntactKV is pre-computed, the pre-filling stage of the quantized LLMs can be accelerated as well. The memory overhead to save IntactKV is also negligible compared with the LLM backbone. For instance, there are only 34 tokens of the system prompt for Vicuna-v1.5-7B, and thus IntactKV takes only 0.13% of the LLM model parameters.
3.2 IntactKV as Trainable Parameters
Since IntactKV is pre-computed and saved offline, it can be treated as extra trainable parameters aside from the LLM backbone to further boost the quantized LLMs. Despite there being no information loss at the pivot tokens, the quantization may still introduce errors to the KV cache during the decoding stage. As shown in Figure 2(a), we calibrate IntactKV to compensate for the quantization error accumulated in the following tokens. While there are various metrics to characterize the quantization discrepancy Frantar et al. (2022); Shao et al. (2024); Liu et al. (2023a), we adopt the mean square error of the transformer layer output between the full-precision LLM and quantized LLM, a simple yet most widely used metric, i.e.,
$\mathcal{L}(\Theta)=\frac{1}{2}\sum_{l=1}^{L}\|f_{l}(\mathbf{w},\mathbf{x})-f_{l}(\hat{\mathbf{w}},\mathbf{x};\Theta)\|_{2}^{2},$ (2)
where $\Theta$ denotes the set of IntactKV, $f_{l}$ is the mapping function for the $l$-th Transformer layer, and $L$ is the number of Transformer layers in LLM. $\mathbf{x}$ is the input sequence, while $\mathbf{w},\hat{\mathbf{w}}$ are full-precision and quantized weights respectively. Note that the full-precision model is only required during the calibration process, and it can be then discarded afterward. It is empirically found that calibration of system prompt IntactKV in SFT models generally gives more improvement than the calibration of [BOS] token in pre-trained LLMs. This matches the intuition that a larger size of IntactKV increases the potential to compensate for quantization errors.
As we focus on the post-training quantization, the training of IntactKV is highly lightweight since the only learnable parameters introduced are IntactKV, i.e., the KV cache of pivot tokens. It takes only as few as 20 epochs on a calibration set with 128 samples. Besides, training with a quantized model further lowers the memory cost.
3.3 Theoretical Analysis
In this section, we provide a theoretical view of how the proposed IntactKV benefits the quantized LLM. For the clarity of presentation, our analysis is built on the self-attention module of a Transformer layer, while it can be readily extended to the FFN module and multiple layers.
Specifically, we denote ${\bm{K}},{\bm{V}}\in\mathbb{R}^{n\times d}$ as the KV cache during the decoding stage, and ${\bm{q}}\in\mathbb{R}^{d}$ is the query vector, where $n$ and $d$ are the sequence length and head dimension. Recall that the output of each attention head ${\bm{h}}\in\mathbb{R}^{d}$ is computed as
${\bm{h}}={\rm softmax}({\bm{q}}{\bm{K}}^{\top}/\sqrt{d}){\bm{V}}{\bm{W}}^{O},$ (3)
where ${\bm{W}}^{O}\in\mathbb{R}^{d\times d}$ is the weight matrix of the projection layer. By quantizing the LLMs, there will be errors accumulated on the KV cache, denoted as $\Delta{\bm{K}},\Delta{\bm{V}}\in\mathbb{R}^{n\times d}$. Therefore, we are interested in showing how $\Delta{\bm{K}}$ and $\Delta{\bm{V}}$ are propagated to the change of attention head $\Delta{\bm{h}}$, and to what extent IntactKV alleviates the distortion.
Theorem 1.
Given the query vector ${\bm{q}}\in\mathbb{R}^{d}$ and the change of KV caches $\Delta{\bm{K}},\Delta{\bm{V}}\in\mathbb{R}^{n\times d}$, the change of the attention head $\Delta{\bm{h}}$ is bounded by
${\|\Delta{\bm{h}}\|}_{2}\leq$ $\,C_{1}{\|\Delta{\bm{K}}\|}_{2,\infty}{\|\Delta{\bm{V}}\|}_{F}\,+$
$+C_{2}{\|\Delta{\bm{K}}\|}_{2,\infty}+C_{3}{\|\Delta{\bm{V}}\|}_{F},$
where $C_{1}=\frac{n^{3/2}}{\sqrt{d}}C_{3}{\|{\bm{q}}\|}_{2},C_{2}=C_{1}{\|{\bm{V}}\|}_{2}$ and $C_{3}={\|{\bm{W}}^{O}\|}_{2}$.
The proof to Theorem 1 can be found in Appendix A. We preserve the terms w.r.t. $\Delta{\bm{K}}$ and $\Delta{\bm{V}}$ of interests, and leave the rest as constants. Note that $\Delta{\bm{K}}$ can be further separated by the pivot tokens $\Delta{\bm{K}}_{p}$ and rest tokens $\𝑝\Delta{\bm{K}}_{\textbackslash p}$, and similar notations hold for $\Delta{\bm{V}}$. Therefore, we have ${\|\Delta{\bm{K}}\|}_{2,\infty}=\max\big{(}{\|\Delta{\bm{K}}_{p}\|}_{2,\infty},{\|\Delta{\bm{K}}_{\textbackslash p}\|}_{2,\infty}\big{)},$ and ${\|\Delta{\bm{V}}\|}_{F}=\sqrt{{\|\Delta{\bm{V}}_{p}\|}_{F}^{2}+{\|\Delta{\bm{V}}_{\textbackslash p}\|}_{F}^{2}}$. With IntactKV  we have ${\|\Delta{\bm{K}}_{p}\|}_{2,\infty}=\|\Delta{\bm{V}}_{p}\|_{F}=0$ since they are generated losslessly, which decreases the upper bound of $\|\Delta{\bm{h}}\|_{2}$. Moreover, it can further reduce the bound by incorporating more pivot tokens. This also aligns with the observation in Figure 2 that a larger size of IntactKV gives a lower MSE of the attention module.
4 Experiments
Method LLaMA-7B LLaMA-13B LLaMA-30B LLaMA-65B LLaMA-2-7B LLaMA-2-13B LLaMA-2-70B
FP16 7.36 6.82 6.15 5.83 7.11 6.58 5.59
RTN 9.15 7.89 6.85 6.33 8.79 7.43 6.12
GPTQ 8.59 7.49 6.73 6.29 9.41 7.25 6.27
OmniQuant 8.24 7.38 6.64 6.17 8.21 7.23 5.96
AWQ 8.26 7.38 6.59 6.16 8.12 7.15 5.91
  +IntactKV${}_{\textrm{[B]}}$ 8.12 7.36 6.54 6.12 8.00 7.12 5.89
Table 1: INT3-group128 quantization results of LLaMA and LLaMA-2 Models on C4 dataset.
Model Method MMLU (0 shot) MMLU (5 shot)
Hums STEM Social Others Avg. Hums STEM Social Others Avg.
Vicuna-v1.5-13B FP16 50.48% 43.70% 62.72% 62.74% 54.54% 51.97% 44.96% 65.26% 62.40% 55.78%
RTN 46.61% 41.32% 58.92% 57.53% 50.69% 47.14% 42.81% 59.38% 58.17% 51.44%
GPTQ 48.35% 40.99% 59.25% 57.99% 51.38% 49.63% 43.04% 60.22% 60.09% 52.95%
OmniQuant 49.52% 41.22% 59.47% 57.74% 51.82% 49.12% 44.40% 60.48% 58.95% 52.86%
AWQ 48.82% 41.72% 61.03% 58.30% 52.16% 49.52% 43.01% 61.72% 58.73% 52.92%
  +IntactKV${}_{\textrm{[B]}}$ 49.31% 42.18% 61.20% 59.28% 52.68% 50.31% 43.37% 61.91% 59.93% 53.58%
Table 2: INT3-group128 quantization results of Vicuna-v1.5-13B on 0-shot and 5-shot MMLU benchmarks.
Task Acc MMLU (5 shot) average Common Sense QA (0 shot) average
Vicuna Family v1.5-7B v1.5-13B v1.3-7B v1.3-13B v1.3-33B v1.5-7B v1.5-13B v1.3-7B v1.3-13B v1.3-33B
FP16 49.84% 55.78% 47.12% 52.10% 59.30% 65.33% 68.38% 64.52% 67.22% 69.53%
RTN 44.62% 51.44% 39.33% 44.56% 53.18% 60.99% 65.40% 59.13% 63.23% 67.19%
GPTQ 43.99% 52.95% 40.12% 47.83% 55.84% 58.09% 65.78% 59.80% 64.03% 66.68%
OmniQuant 46.54% 52.86% 43.18% 47.92% 55.12% 61.73% 65.23% 61.08% 64.81% 67.58%
AWQ 46.45% 52.92% 43.08% 48.56% 56.09% 61.86% 66.01% 60.94% 64.53% 67.67%
  +IntactKV${}_{\textrm{[B]}}$ 46.87% 53.58% 44.67% 49.05% 56.91% 62.49% 66.93% 61.93% 65.02% 67.90%
Table 3: INT3-group128 quantization results of various Vicuna models on 5-shot MMLU and 0-shot QA tasks.
4.1 Settings
Models.
We evaluate the proposed IntactKV on various sizes of open-sourced LLMs, including LLaMA (Touvron et al., 2023a) (7B-65B), LLaMA-2 (Touvron et al., 2023b) (7B-70B), Vicuna-v1.3 (7B-33B) and Vicuna-v1.5 (Touvron et al., 2023b) (7B-13B). We denote models that keep intact [BOS] KV as IntactKV${}_{\textrm{[B]}}$, and models that keep intact system prompt KV as IntactKV${}_{\textrm{[P]}}$.
Quantization Methods.
We mainly consider weight-only quantization methods, including round-to-nearest quantization (RTN), GPTQ (Frantar et al., 2022), the state-of-the-art OmniQuant (Shao et al., 2024) and AWQ (Lin et al., 2023). For GPTQ, we use AutoGPTQ [^1] https://github.com/AutoGPTQ/AutoGPTQ with C4 calibration set following GPTQ paper (Frantar et al., 2022) to reproduce all results. For AWQ [^2] https://github.com/mit-han-lab/llm-awq and OmniQuant [^3] https://github.com/OpenGVLab/OmniQuant, we directly load the officially released quantization parameters of LLaMA models for evaluation and reproduce results on Vicuna models with their official code. We use Pile (Gao et al., 2020) calibration set for AWQ and WikiText2 (Merity et al., 2016) calibration set for OmniQuant, following (Lin et al., 2023; Shao et al., 2024). We adopt asymmetric group-wise quantization with a group size of 128 and mainly focus on INT3 and INT4 quantization since INT8 is empirically lossless on various task metrics (Dettmers et al., 2022).
While IntactKV can be readily combined with these existing quantization methods, we mainly apply IntactKV for AWQ in the main experiments due to the space limitation and the state-of-the-art performance of AWQ. We shall provide a more comprehensive evaluation of combining IntactKV with other quantization methods in Section 4.3. Moreover, aside from weight-only quantization, the proposed IntactKV can be similarly applied for KV cache quantization, as detailed in Section 4.4. Nonetheless, it is non-trivial to apply IntactKV for activation quantization, since the quantization over KV cache cannot keep them intact. We leave it as future work and more discussions are provided in Section 6.
Evaluation.
For pre-trained LLMs (i.e., LLaMA and LLaMA-2), we report the perplexity (PPL) of language generation on C4 (Raffel et al., 2020) and WikiText2 (Merity et al., 2016) dataset. For SFT models (i.e., Vicuna-v1.3 and v1.5), we conduct evaluation over a wide range of downstream tasks. We test the zero and five-shot performance on the Massively Multitask Language Understanding (MMLU) (Hendrycks et al., 2020) benchmark. Meanwhile, we also evaluate seven zero-shot commonsense QA tasks: OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2021), ARC-Challenge, ARC-Easy (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), and LAMBADA (Paperno et al., 2016). Additionally, we evaluate quantized Vicuna on MT-bench (Zheng et al., 2023), a high-quality dataset consisting of 80 open-ended multi-turn questions, to gauge their alignment with human preferences. The responses generated by quantized models are judged by GPT-4 with a total score of 10. More evaluation details can be found in Appendix E.
Implementation Details
Model #bits Method OBQA WinoGrande ARC-C ARC-E BoolQ HellaSwag LAMBADA Avg
Vicuna-v1.5-13B FP16 - 45.40% 71.51% 50.68% 74.87% 85.29% 77.50% 73.43% 68.38%
w3g128 RTN 42.00% 70.01% 47.44% 72.77% 82.20% 74.18% 69.20% 65.40%
GPTQ 42.40% 69.53% 48.46% 71.84% 83.76% 74.48% 70.00% 65.78%
OmniQuant 43.40% 68.51% 47.53% 71.09% 82.32% 73.92% 69.84% 65.23%
AWQ 44.00% 68.75% 48.21% 71.76% 83.58% 75.09% 70.68% 66.01%
  +IntactKV${}_{\textrm{[B]}}$ 45.40% 70.32% 48.38% 72.14% 85.20% 75.23% 71.86% 66.93%
w4g128 RTN 44.20% 70.80% 48.98% 73.82% 84.68% 76.36% 73.04% 67.41%
GPTQ 45.80% 70.96% 50.51% 73.99% 85.47% 76.70% 73.43% 68.12%
OmniQuant 43.80% 70.24% 49.74% 73.61% 84.59% 76.35% 72.54% 67.27%
AWQ 44.00% 72.06% 49.15% 73.44% 85.17% 77.00% 72.77% 67.66%
  +IntactKV${}_{\textrm{[B]}}$ 45.40% 73.09% 49.57% 74.45% 85.66% 77.32% 72.75% 68.32%
Table 4: Quantization results of Vicuna-v1.5-13B on seven 0-shot commonsense QA tasks.
For evaluation on PPL, MMLU, and commonsense QA tasks, we adopt IntactKV${}_{\textrm{[B]}}$ that only includes [BOS] KV since the input sequence of these tasks does not use any system prompt. For evaluation of SFT models on MT-bench, we adopt IntactKV${}_{\textrm{[P]}}$ to keep an intact system prompt KV cache. The system prompt of Vicuna can be found in Appendix B. For training the cached IntactKV, we randomly sample 128 samples from ShareGPT [^4] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered dataset as our calibration dataset, consisting of multi-turn ChatGPT (OpenAI, 2022) conversations. The layer-wise MSE defined in Equation 2 is calculated on the response of ChatGPT. We use AdamW optimizer with learning rate $2\times 10^{-4}$, training for 160 optimizer update steps with a gradient accumulation step of 16, i.e., 20 epochs. As mentioned in Section 3.2, training IntactKV${}_{\textrm{[B]}}$ leads to comparable performance compared with vanilla IntactKV. Instead, the calibration of IntactKV${}_{\textrm{[P]}}$ has more potential to improve quantized LLMs with longer system prompt. Thus, we primarily evaluate the trainable IntactKV${}_{\textrm{[P]}}$ with system prompt as pivot tokens in the following experiments.
4.2 Main Results
Results on Language Generation Tasks.
We first integrate our proposed IntactKV with AWQ on LLaMA and LLaMA-2 models. The effect of this integration on model performance was measured by the perplexity (PPL) metric, with results on the C4 dataset detailed in Table 1, and results on the WikiText2 dataset in Appendix F.1. As indicated in these tables, IntactKV notably enhances the generative capabilities of AWQ-quantized models across various sizes, consistently surpassing the prior state-of-the-art (SOTA) method, OmniQuant. These findings demonstrate the efficacy of IntactKV in improving quantized LLMs and particularly highlight the effectiveness of utilizing the KV cache from full-precision models.
Results on MMLU Tasks.
#bits Method Vicuna-v1.5-7B Vicuna-v1.5-13B
FP16 - 5.31 5.52
w3g128 RTN 4.34 5.13
OmniQuant 4.78 5.05
AWQ 4.74 5.17
  +IntactKV${}_{\textrm{[P]}}$ 4.68 5.34
  +IntactKV${}_{\textrm{[P]}}$+FT 4.84 5.44
w4g128 RTN 5.18 5.47
OmniQuant 5.09 5.48
AWQ 5.22 5.28
  +IntactKV${}_{\textrm{[P]}}$ 5.32 5.35
  +IntactKV${}_{\textrm{[P]}}$+FT 5.36 5.50
Table 5: GPT-4 evaluation of quantized Vicuna-v1.5 models on MT-Bench. The scores are on a scale of 10.
For SFT models, we implement IntactKV on the AWQ-quantized Vicuna models and evaluate the multi-task problem-solving ability on the MMLU benchmark. Table 2 presents the detailed zero-shot and five-shot results for Vicuna-v1.5-13B. The results demonstrate that IntactKV significantly enhances the performance of the AWQ-quantized model across all categories of tasks for Vicuna-v1.5-13B. Moreover, the performance of other model sizes under the five-shot setting is outlined in Table 3. Remarkably, AWQ+IntactKV exhibits superior performance over OmniQuant, achieving an average improvement of 1.09% across five model sizes. More results on MMLU are provided in Appendix F.2.
Results on Commonsense QA Tasks.
#bits Method Vicuna-v1.5-7B Vicuna-v1.5-13B
FP16 - 5.31 5.52
w3g128 RTN 4.34 5.13
  +IntactKV${}_{\textrm{[P]}}$ 4.72 5.27
  +IntactKV${}_{\textrm{[P]}}$+FT 4.73 5.30
OmniQuant 4.78 5.05
  +IntactKV${}_{\textrm{[P]}}$ 4.94 5.10
  +IntactKV${}_{\textrm{[P]}}$+FT 4.85 5.24
AWQ 4.74 5.17
  +IntactKV${}_{\textrm{[P]}}$ 4.68 5.34
  +IntactKV${}_{\textrm{[P]}}$+FT 4.84 5.44
Table 6: Compatibility of IntactKV with various quantization methods. The scores are on a scale of 10.
Figure 4: Results of KV cache quantization with different bit-widths on 5-shot MMLU benchmark. Note that this is additional to INT3/4 weight quantization. Blue and red lines indicate quantizing model weights to INT3 and INT4, respectively. More experiment details can be found in Appendix G.
We further evaluate the quantized Vicuna models on zero-shot commonsense QA tasks. The results of Vicuna-v1.5-13B, as detailed in Table 4, indicate that IntactKV enables significant improvements over AWQ. Notably, our AWQ+IntactKV largely surpasses GPTQ and OmniQuant, especially under low-bit quantization. Additionally, Table 3 presents the average accuracy for various sizes of Vicuna models. In these evaluations, our AWQ+IntactKV consistently achieves the best zero-shot performance, which strongly demonstrates the efficacy of our proposed IntactKV. More results on Commonsense QA tasks can be found in Appendix F.3.
Results on MT-Bench.
To evaluate the quantized models’ generation capabilities in multi-turn conversations and their alignment with human preferences, we use GPT-4 to score the responses of quantized models on MT-Bench. We also fine-tune IntactKV, denoted as IntactKV+FT. From Table 5, IntactKV+FT significantly boosts the quantized model and consistently surpasses the GPTQ and OmniQuant for both INT3-g128 and INT4-g128 settings. For example, the 3-bit Vicuna-v1.5-13B quantized by AWQ has been improved from 5.17 to 5.34 by using the IntactKV, which can be further boosted to 5.44 with further fine-tuning. Remarkably, with trainable IntactKV, the setting of INT4-g128 even matches the full-precision model, while all other methods clearly lag behind the full-precision model. These results demonstrate the effectiveness of treating IntactKV as trainable parameters. Notably, the training process for the 7B model takes only 10 minutes on a single computing device, which is quite lightweight.
4.3 IntactKV with Other PTQ Methods
To assess IntactKV’s compatibility with different quantization methods, we build IntactKV upon various PTQ methods (i.e., RTN, OmniQuant, and AWQ) and evaluate on MT-bench with INT3-g128 quantization. From Table 6, IntactKV with different PTQ methods lead to an average boost of 0.16 in the final score for Vicuna-v1.5-7B, and 0.12 for Vicuna-v1.5-13B. Fine-tuning IntactKV further improves the performance of quantized models. Our IntactKV+FT consistently surpasses the performance of original quantized models for all three different quantization methods, raising the final score by 0.19 for Vicuna-v1.5-7B and 0.21 for Vicuna-v1.5-13B on average. These findings validate IntactKV’s capability of acting as a lightweight plugin, consistently improving various quantized models with negligible extra costs.
4.4 Extension to KV Cache Quantization
IntactKV can be readily applied to KV cache quantization to further decrease memory requirements. Similar to weight-only quantization, the quantized KV cache needs to be de-quantized before the matrix multiplication. Nonetheless, the proposed IntactKV is kept in FP16 since it is more sensitive to quantization. Also, the concatenation between the FP16 IntactKV and the rest de-quantized KV cache incurs negligible extra overhead. From Figure 4, IntactKV notably improves AWQ across different models and KV cache bit widths under the INT3 weight quantization. For INT4 weight quantization, IntactKV still gains an average accuracy increase of 0.27% over AWQ. We also notice that quantizing the KV cache to INT8 leads to almost no performance drop on the MMLU benchmark. When equipped with IntactKV, INT8 KV cache can even surpass vanilla AWQ-quantized models with FP16 KV cache, especially under INT3 weight quantization.
5 Conclusions
In this paper, we propose IntactKV, a simple and easy-to-combine method to improve large language model quantization. The research is motivated by the previously overlooked outliers over pivot tokens, which lead to attention sinks that are critical to the performance of quantized LLMs. By generating IntactKV with the full-precision model, the quantization error accumulated over the attention scores can be effectively alleviated. IntactKV can also be calibrated as additional parameters to the LLM backbone, further improving the quantized LLMs. Experiments show that combining the proposed IntactKV gives consistent improvement on various sizes of LLMs and across multiple downstream tasks, leading to new state-of-the-art results for large language model quantization.
6 Limitations
Firstly, it is non-trivial to integrate IntactKV into activation quantization. For activation quantization, the whole KV cache needs to be quantized to low bits to exploit integer multiplications in self-attention, which contradicts our idea of keeping pivot tokens’ KV cache intact. Since IntactKV is treated as extra cached parameters, a straightforward solution is to quantize IntactKV with suitable quantization methods such as GPTQ. We leave this as future work. Secondly, more experiments may be needed for LLM evaluation. LLMs are being applied to a wide range of tasks, posing high demands on various model abilities. When quantizing LLMs to low bits, these abilities may be affected to varying degrees. Therefore, a comprehensive evaluation is required to gauge the capabilities of quantized LLMs. Although we experiment on several downstream tasks, such as PPL, MMLU, commonsense QA, and MT-bench, we note that this may not be enough to assess all abilities of LLMs. For example, how long context affects quantized models still remains unknown.
7 Ethics Statement
The development of LLM quantization techniques can further democratize LLMs, lowering the costs of LLM serving and enabling more people to get access to advanced AI assistants. Nonetheless, LLM itself may inherit certain social biases from training data concerning gender, race, etc. Quantization can not mitigate such biases. Therefore, caution must be taken when using quantized LLMs.
Appendix A Proof of Theorem 1
Method LLaMA-7B LLaMA-13B LLaMA-30B LLaMA-65B LLaMA-2-7B LLaMA-2-13B LLaMA-2-70B
FP16 5.69 5.08 4.09 3.52 5.12 4.57 3.12
RTN 6.98 5.88 4.84 4.22 6.21 5.17 3.74
GPTQ 6.62 5.68 4.75 4.20 6.68 5.14 3.79
OmniQuant 6.18 5.46 4.57 4.00 5.77 4.96 3.58
AWQ 6.34 5.53 4.60 3.95 5.82 4.97 3.53
  +IntactKV${}_{\textrm{[B]}}$ 6.23 5.49 4.54 3.89 5.73 4.95 3.50
Table 7: INT3-group128 quantization results of LLaMA and LLaMA-2 Models on WikiText2 dataset.
Model Method MMLU (0 shot) MMLU (5 shot)
Hums STEM Social Others Avg. Hums STEM Social Others Avg.
Vicuna-v1.5-7B FP16 45.40% 38.67% 56.16% 55.92% 48.74% 45.78% 39.50% 58.14% 57.46% 49.84%
RTN 42.06% 34.16% 50.47% 50.59% 44.17% 40.68% 38.60% 50.31% 50.56% 44.62%
GPTQ 39.89% 33.00% 48.10% 48.46% 42.19% 40.30% 36.28% 50.76% 50.09% 43.99%
OmniQuant 42.57% 35.98% 51.64% 53.55% 45.68% 42.95% 37.57% 53.59% 53.39% 46.54%
AWQ 42.08% 35.55% 51.61% 51.54% 44.95% 42.55% 38.93% 53.10% 52.78% 46.45%
  +IntactKV${}_{\textrm{[B]}}$ 42.42% 35.42% 51.71% 51.57% 45.06% 42.95% 38.60% 54.37% 53.15% 46.87%
Vicuna-v1.5-13B FP16 50.48% 43.70% 62.72% 62.74% 54.54% 51.97% 44.96% 65.26% 62.40% 55.78%
RTN 46.61% 41.32% 58.92% 57.53% 50.69% 47.14% 42.81% 59.38% 58.17% 51.44%
GPTQ 48.35% 40.99% 59.25% 57.99% 51.38% 49.63% 43.04% 60.22% 60.09% 52.95%
OmniQuant 49.52% 41.22% 59.47% 57.74% 51.82% 49.12% 44.40% 60.48% 58.95% 52.86%
AWQ 48.82% 41.72% 61.03% 58.30% 52.16% 49.52% 43.01% 61.72% 58.73% 52.92%
  +IntactKV${}_{\textrm{[B]}}$ 49.31% 42.18% 61.20% 59.28% 52.68% 50.31% 43.37% 61.91% 59.93% 53.58%
Vicuna-v1.3-7B FP16 44.31% 36.28% 53.23% 53.70% 46.71% 44.23% 38.34% 53.82% 53.15% 47.12%
RTN 38.09% 31.58% 42.35% 44.32% 39.06% 36.81% 32.77% 43.87% 44.79% 39.33%
GPTQ 39.09% 32.57% 44.59% 46.73% 40.66% 36.94% 33.90% 45.08% 45.81% 40.12%
OmniQuant 41.36% 33.96% 47.42% 47.69% 42.56% 41.15% 35.45% 48.20% 48.58% 43.18%
AWQ 40.49% 32.44% 47.06% 49.57% 42.29% 39.64% 36.22% 48.72% 49.11% 43.08%
  +IntactKV${}_{\textrm{[B]}}$ 41.76% 32.94% 47.74% 49.72% 43.01% 41.93% 36.58% 50.37% 50.77% 44.67%
Vicuna-v1.3-13B FP16 47.89% 39.96% 58.86% 57.34% 50.77% 49.78% 40.46% 60.61% 58.24% 52.10%
RTN 42.06% 32.87% 47.61% 49.51% 43.02% 42.42% 34.46% 50.34% 51.57% 44.56%
GPTQ 45.06% 35.88% 52.23% 51.26% 46.09% 45.82% 37.57% 54.83% 53.64% 47.83%
OmniQuant 43.29% 36.65% 51.64% 53.05% 45.95% 45.29% 37.84% 55.02% 54.38% 47.92%
AWQ 45.14% 36.18% 52.55% 53.79% 46.84% 46.65% 37.64% 55.54% 54.87% 48.56%
  +IntactKV${}_{\textrm{[B]}}$ 45.91% 36.65% 53.75% 54.60% 47.64% 46.57% 38.40% 56.03% 55.95% 49.05%
Vicuna-v1.3-33B FP16 53.73% 44.14% 67.63% 63.54% 56.98% 57.66% 46.32% 69.32% 64.25% 59.30%
RTN 49.88% 40.13% 61.33% 58.42% 52.26% 51.26% 42.54% 61.75% 57.71% 53.18%
GPTQ 51.22% 40.03% 61.85% 59.47% 53.05% 54.05% 44.04% 64.35% 61.35% 55.84%
OmniQuant 51.14% 42.08% 63.60% 59.84% 53.93% 53.77% 43.80% 63.47% 59.69% 55.12%
AWQ 51.69% 42.74% 63.41% 61.38% 54.57% 54.56% 44.10% 65.36% 60.67% 56.09%
  +IntactKV${}_{\textrm{[B]}}$ 52.09% 42.68% 63.70% 62.03% 54.91% 55.79% 44.90% 65.62% 61.47% 56.91%
Table 8: INT3-group128 quantization results of Vicuna models on 0-shot and 5-shot MMLU benchmarks.
Model Method MMLU (0 shot) MMLU (5 shot)
Hums STEM Social Others Avg. Hums STEM Social Others Avg.
Vicuna-v1.5-7B FP16 45.40% 38.67% 56.16% 55.92% 48.74% 45.78% 39.50% 58.14% 57.46% 49.84%
RTN 44.65% 38.47% 53.95% 54.41% 47.61% 44.87% 39.13% 56.45% 55.34% 48.59%
GPTQ 44.87% 37.08% 54.44% 53.86% 47.37% 45.44% 38.83% 57.33% 56.14% 49.10%
OmniQuant 45.61% 38.70% 55.64% 56.23% 48.78% 45.21% 38.77% 57.17% 57.37% 49.25%
AWQ 45.08% 37.41% 55.64% 55.31% 48.11% 45.44% 38.97% 56.94% 55.74% 48.95%
  +IntactKV${}_{\textrm{[B]}}$ 45.25% 37.51% 55.93% 55.58% 48.31% 45.33% 39.60% 57.36% 55.74% 49.14%
Vicuna-v1.5-13B FP16 50.48% 43.70% 62.72% 62.74% 54.54% 51.97% 44.96% 65.26% 62.40% 55.78%
RTN 50.01% 43.41% 62.33% 62.00% 54.06% 51.31% 43.14% 63.54% 61.63% 54.61%
GPTQ 50.20% 42.31% 61.62% 61.41% 53.60% 50.10% 43.97% 62.72% 61.01% 54.07%
OmniQuant 50.01% 43.70% 62.33% 62.00% 54.12% 51.67% 43.87% 63.34% 61.81% 54.89%
AWQ 50.10% 42.94% 61.68% 61.66% 53.77% 52.31% 44.43% 63.18% 61.84% 55.20%
  +IntactKV${}_{\textrm{[B]}}$ 50.14% 42.84% 61.78% 61.91% 53.84% 52.31% 44.37% 63.67% 61.91% 55.31%
Vicuna-v1.3-7B FP16 44.31% 36.28% 53.23% 53.70% 46.71% 44.23% 38.34% 53.82% 53.15% 47.12%
RTN 42.78% 36.55% 51.74% 51.48% 45.41% 42.23% 37.08% 52.10% 51.94% 45.53%
GPTQ 43.40% 34.46% 52.06% 53.45% 45.70% 43.78% 36.41% 53.49% 52.41% 46.32%
OmniQuant 42.81% 34.72% 52.26% 52.00% 45.26% 43.21% 37.81% 52.62% 53.24% 46.43%
AWQ 43.53% 36.22% 53.01% 52.53% 46.11% 43.36% 37.74% 53.46% 52.68% 46.52%
  +IntactKV${}_{\textrm{[B]}}$ 43.57% 36.51% 52.29% 53.27% 46.20% 43.51% 37.44% 53.17% 52.62% 46.43%
Vicuna-v1.3-13B FP16 47.89% 39.96% 58.86% 57.34% 50.77% 49.78% 40.46% 60.61% 58.24% 52.10%
RTN 47.16% 39.00% 56.52% 56.63% 49.64% 49.25% 39.63% 57.85% 57.74% 51.03%
GPTQ 46.95% 39.30% 57.39% 56.23% 49.74% 49.05% 39.46% 59.02% 57.65% 51.16%
OmniQuant 47.93% 39.50% 57.98% 57.31% 50.48% 49.18% 39.86% 59.41% 58.14% 51.49%
AWQ 48.03% 39.43% 56.94% 56.76% 50.15% 49.44% 40.49% 59.57% 57.65% 51.63%
  +IntactKV${}_{\textrm{[B]}}$ 47.91% 39.60% 57.69% 56.79% 50.31% 49.54% 40.23% 60.12% 57.71% 51.74%
Vicuna-v1.3-33B FP16 53.73% 44.14% 67.63% 63.54% 56.98% 57.66% 46.32% 69.32% 64.25% 59.30%
RTN 53.18% 44.27% 66.88% 62.95% 56.52% 56.73% 45.73% 68.09% 62.49% 58.18%
GPTQ 52.92% 44.90% 67.05% 63.66% 56.77% 57.13% 45.96% 67.63% 63.11% 58.41%
OmniQuant 53.22% 44.73% 67.53% 63.05% 56.80% 56.68% 45.46% 68.67% 62.49% 58.24%
AWQ 53.22% 44.40% 67.63% 63.54% 56.87% 56.85% 45.69% 68.80% 63.66% 58.65%
  +IntactKV${}_{\textrm{[B]}}$ 53.37% 44.40% 67.50% 63.63% 56.91% 57.07% 45.96% 68.51% 63.63% 58.70%
Table 9: INT4-group128 quantization results of Vicuna models on 0-shot and 5-shot MMLU benchmarks.
Model #bits Method OBQA WinoGrande ARC-C ARC-E BoolQ HellaSwag LAMBADA Avg
Vicuna-V1.5-7B FP16 - 45.00% 69.53% 45.73% 71.25% 80.92% 73.78% 71.12% 65.33%
w3g128 RTN 39.40% 66.22% 43.34% 67.72% 77.40% 71.71% 61.17% 60.99%
GPTQ 37.80% 63.30% 41.81% 64.73% 74.46% 65.97% 58.57% 58.09%
OmniQuant 41.40% 66.22% 43.52% 67.26% 77.13% 70.35% 66.23% 61.73%
AWQ 40.80% 67.01% 43.34% 67.55% 78.32% 71.03% 64.97% 61.86%
  +IntactKV${}_{\textrm{[B]}}$ 42.20% 67.64% 41.98% 68.52% 79.02% 71.24% 66.82% 62.49%
w4g128 RTN 43.80% 68.82% 45.39% 70.33% 81.10% 73.31% 69.09% 64.55%
GPTQ 44.40% 69.93% 44.71% 70.45% 80.76% 73.72% 70.29% 64.89%
OmniQuant 43.00% 68.75% 44.28% 70.50% 81.01% 72.82% 70.10% 64.35%
AWQ 43.40% 68.75% 45.39% 70.66% 81.35% 73.43% 69.47% 64.64%
  +IntactKV${}_{\textrm{[B]}}$ 44.00% 68.90% 45.90% 71.63% 82.29% 73.52% 69.61% 65.12%
Vicuna-v1.5-13B FP16 - 45.40% 71.51% 50.68% 74.87% 85.29% 77.50% 73.43% 68.38%
w3g128 RTN 42.00% 70.01% 47.44% 72.77% 82.20% 74.18% 69.20% 65.40%
GPTQ 42.40% 69.53% 48.46% 71.84% 83.76% 74.48% 70.00% 65.78%
OmniQuant 43.40% 68.51% 47.53% 71.09% 82.32% 73.92% 69.84% 65.23%
AWQ 44.00% 68.75% 48.21% 71.76% 83.58% 75.09% 70.68% 66.01%
  +IntactKV${}_{\textrm{[B]}}$ 45.40% 70.32% 48.38% 72.14% 85.20% 75.23% 71.86% 66.93%
w4g128 RTN 44.20% 70.80% 48.98% 73.82% 84.68% 76.36% 73.04% 67.41%
GPTQ 45.80% 70.96% 50.51% 73.99% 85.47% 76.70% 73.43% 68.12%
OmniQuant 43.80% 70.24% 49.74% 73.61% 84.59% 76.35% 72.54% 67.27%
AWQ 44.00% 72.06% 49.15% 73.44% 85.17% 77.00% 72.77% 67.66%
  +IntactKV${}_{\textrm{[B]}}$ 45.40% 73.09% 49.57% 74.45% 85.66% 77.32% 72.75% 68.32%
Vicuna-V1.3-7B FP16 - 43.80% 69.46% 44.54% 71.89% 78.07% 73.93% 69.98% 64.52%
w3g128 RTN 41.00% 64.64% 39.08% 65.03% 75.81% 68.59% 59.73% 59.13%
GPTQ 39.60% 64.17% 42.24% 65.95% 72.63% 70.36% 63.65% 59.80%
OmniQuant 42.00% 67.88% 40.02% 66.75% 75.41% 70.52% 64.97% 61.08%
AWQ 41.60% 67.96% 38.65% 65.95% 76.36% 71.27% 64.76% 60.94%
  +IntactKV${}_{\textrm{[B]}}$ 43.60% 68.43% 39.16% 67.30% 77.28% 71.20% 66.54% 61.93%
w4g128 RTN 41.80% 67.80% 44.45% 69.91% 75.57% 73.25% 67.26% 62.86%
GPTQ 44.20% 68.75% 43.60% 70.88% 75.20% 73.39% 69.42% 63.63%
OmniQuant 42.40% 68.98% 44.03% 71.04% 76.97% 73.19% 69.49% 63.73%
AWQ 42.80% 67.09% 43.43% 71.34% 76.36% 73.46% 68.78% 63.32%
  +IntactKV${}_{\textrm{[B]}}$ 43.80% 68.59% 42.92% 71.84% 76.79% 73.49% 69.57% 63.86%
Vicuna-V1.3-13B FP16 - 45.40% 71.03% 47.70% 73.70% 82.81% 77.00% 72.91% 67.22%
w3g128 RTN 42.20% 69.77% 44.03% 68.14% 80.52% 73.28% 64.68% 63.23%
GPTQ 42.00% 68.35% 44.88% 70.08% 80.67% 74.24% 68.00% 64.03%
OmniQuant 43.60% 70.56% 44.62% 70.66% 81.28% 74.13% 68.79% 64.81%
AWQ 42.60% 68.75% 45.90% 69.57% 80.80% 74.78% 69.30% 64.53%
  +IntactKV${}_{\textrm{[B]}}$ 43.20% 69.46% 46.16% 69.74% 81.80% 75.11% 69.67% 65.02%
w4g128 RTN 43.80% 71.19% 47.01% 72.64% 82.35% 76.19% 71.22% 66.34%
GPTQ 44.60% 70.01% 47.87% 73.32% 82.23% 76.55% 71.78% 66.62%
OmniQuant 45.80% 71.35% 46.08% 72.14% 82.35% 76.34% 71.38% 66.49%
AWQ 44.00% 70.01% 46.67% 72.64% 82.66% 76.43% 71.74% 66.31%
  +IntactKV${}_{\textrm{[B]}}$ 45.60% 71.19% 47.10% 73.32% 82.72% 76.95% 71.38% 66.89%
Vicuna-V1.3-33B FP16 - 47.80% 74.35% 51.79% 74.71% 83.91% 80.38% 73.74% 69.53%
w3g128 RTN 46.00% 72.45% 49.83% 71.63% 82.75% 77.86% 69.82% 67.19%
GPTQ 44.40% 72.14% 47.61% 69.65% 83.49% 77.60% 71.86% 66.68%
OmniQuant 45.60% 73.32% 47.95% 72.52% 83.58% 78.05% 72.06% 67.58%
AWQ 45.40% 72.77% 50.77% 72.56% 82.42% 78.33% 71.45% 67.67%
  +IntactKV${}_{\textrm{[B]}}$ 44.80% 73.56% 51.11% 72.60% 82.78% 78.55% 71.90% 67.90%
w4g128 RTN 46.60% 73.95% 51.96% 74.07% 83.43% 80.01% 73.43% 69.06%
GPTQ 47.00% 73.48% 50.85% 73.06% 83.67% 80.31% 72.50% 68.70%
OmniQuant 47.60% 73.72% 51.96% 73.32% 83.58% 79.85% 73.57% 69.09%
AWQ 46.20% 73.40% 50.60% 73.53% 84.04% 79.91% 73.39% 68.72%
  +IntactKV${}_{\textrm{[B]}}$ 45.60% 73.24% 50.94% 74.12% 84.28% 79.70% 73.14% 68.72%
Table 10: Quantization results of Vicuna models on seven 0-shot commonsense QA tasks.
Model Download URL
LLaMA-2-7B https://huggingface.co/meta-llama/Llama-2-7b
LLaMA-2-13B https://huggingface.co/meta-llama/Llama-2-13b
LLaMA-2-70B https://huggingface.co/meta-llama/Llama-2-70b
Vicuna-v1.3-7B https://huggingface.co/lmsys/vicuna-7b-v1.3
Vicuna-v1.3-13B https://huggingface.co/lmsys/vicuna-13b-v1.3
Vicuna-v1.3-33B https://huggingface.co/lmsys/vicuna-33b-v1.3
Vicuna-v1.5-7B https://huggingface.co/lmsys/vicuna-7b-v1.5
Vicuna-v1.5-13B https://huggingface.co/lmsys/vicuna-13b-v1.5
Table 11: Download links to officially released LLMs.
Proof.
Denote the output of the softmax function as the score ${\bm{s}}$, i.e., ${\bm{s}}={\rm softmax}(\frac{{\bm{q}}{\bm{K}}^{\top}}{\sqrt{d}})$, and also define the error output from the softmax function as $\Delta{\bm{s}}$. To show the error of the attention head, we first justify how the error propagates from the score to the attention head.
${\|\Delta{\bm{h}}\|}_{2}=\left\|[({\bm{s}}+\Delta{\bm{s}})({\bm{V}}+\Delta{\bm{V}})-{\bm{s}}{\bm{V}}]{\bm{W}}^{O}\right\|_{2}$
$\leq\left({\|\Delta{\bm{s}}\|}_{2}{\|{\bm{V}}+\Delta{\bm{V}}\|}_{2}+{\|{\bm{s}}\|}_{2}{\|\Delta{\bm{V}}\|}_{2}\right){\|{\bm{W}}^{O}\|}_{2}$
$\leq\big{(}{\|\Delta{\bm{s}}\|}_{2}({\|{\bm{V}}\|}_{2}\!+\!{\|\Delta{\bm{V}}\|}_{F})+{\|\Delta{\bm{V}}\|}_{F}\big{)}{\|{\bm{W}}^{O}\|}_{2},$
where the inequalities are because
${\|{\bm{x}}+{\bm{y}}\|}_{2}\leq{\|{\bm{x}}\|}_{2}\!+\!{\|{\bm{y}}\|}_{2},~{}~{}{\|{\bm{s}}{\bm{V}}\|}_{2}\leq{\|{\bm{s}}\|}_{2}{\|{\bm{V}}\|}_{2},$
and ${\|{\bm{s}}\|}_{2}\leq{\|{\bm{s}}\|}_{1}=1,{\|{\bm{V}}\|}_{2}\leq{\|{\bm{V}}\|}_{F}$.
Next, we characterize the error of score ${\|\Delta{\bm{s}}\|}_{2}$. This is not easy as the error propagates through the softmax function. To proceed, we need the relative condition number of the softmax function. As indicated in Blanchard et al. (2021),
$\!\!\frac{{\|{\rm softmax}({\bm{x}}\!+\!\Delta{\bm{x}})\!-\!{\rm softmax}({\bm{x}})\|}_{\infty}}{{\|{\rm softmax}({\bm{x}})\|}_{\infty}}\!\!\leq\!\kappa({\bm{x}})\frac{{\|\Delta{\bm{x}}\|}_{\infty}}{{\|{\bm{x}}\|}_{\infty}}\!,$
where $\kappa({\bm{x}})=n{\|{\bm{x}}\|}_{\infty}$ (${\bm{x}}\in\mathbb{R}^{n}$) is an upper bound of the relative condition number of the softmax function. Let ${\bm{x}}={\bm{q}}{\bm{K}}^{\top}/\sqrt{d}$ and $\Delta{\bm{x}}={\bm{q}}\Delta{\bm{K}}^{\top}/\sqrt{d}$, we have
$\frac{{\|\Delta{\bm{s}}\|}_{\infty}}{{\|{\bm{s}}\|}_{\infty}}\leq n{\|\Delta{\bm{x}}\|}_{\infty}\leq\frac{n}{\sqrt{d}}{\|{\bm{q}}\|}_{2}{\|\Delta{\bm{K}}\|}_{2,\infty}.$
Considering that the output of the softmax function is a probability, we have ${\|{\bm{s}}\|}_{\infty}\leq 1$. Therefore, we obtain
${\|\Delta{\bm{s}}\|}_{2}\leq\sqrt{n}{\|\Delta{\bm{s}}\|}_{\infty}\leq\frac{n^{2/3}}{\sqrt{d}}{\|{\bm{q}}\|}_{2}{\|\Delta{\bm{K}}\|}_{2,\infty}.$
Combining the above ingredients, we derive the main results of the Theorem 1.
Appendix B System Prompt of Vicuna Models
Figure 5: System Prompt of Vicuna Models.
Appendix C Visualization of Activations and Attention Map
C.1 Implementation Details
We use ShareGPT dataset for our visualizations, where each sample starts with Vicuna system prompt of length 34. We use a randomly sampled sequence of length 128 to visualize the output activations and plot the corresponding attention map of the first 64 tokens. The attention score is mean pooled over different heads.
C.2 Visualization of LLaMA Models
We provide more visualizations of the output activations and attention map of LLaMA models in Figure. 612. Similar to our observations in Section 2, we find that pivot tokens only appear at the very beginning of the input sequence, and [BOS] always serves as a pivot token.
C.3 Visualization of Vicuna Models
We provide more visualizations of the output activations and attention map of Vicuna models in Figure. 1317. Although Vicuna models demonstrate stronger performance than LLaMA models of the same size, we are surprised to find that the position of pivot tokens remains unchanged for Vicuna and LLaMA models of the same size. Besides, as shown in Figure. 1317, we find that the Vicuna system prompt is enough to cover all the pivot tokens in all Vicuna models.
Appendix D Experiment Details of Figure 2
We plot the quantization loss of the last Transformer layer as well as the total quantization loss of all attention layers with respect to the size of IntactKV on four different models, i.e., LLaMA-13B, LLaMA-30B, LLaMA-2-7B, and LLaMA-2-70B, covering different model types and model sizes. We use lossless IntactKV generated by the full-precision model to quantify the effect of IntactKV on the quantized model. IntactKV of size $s$ can ensure that the KV cache of the first $s$ tokens of the input sequence is generated by the full-precision model and thus lossless. Quantization loss is computed with MSE loss between the output activations of the quantized model and the full-precision model. We sample 128 sequences from the ShareGPT dataset to construct the validation set, each with a common prompt prefix of length 34. MSE loss is calculated on the tokens after the common prompt prefix. We quantize the model weights to 3 bits using round-to-nearest quantization with a group size of 128.
Appendix E Evaluation Details
PPL.
We evaluate PPL following the new evaluation setting in GPTQ official code [^5] https://github.com/ist-daslab/gptq, except that we substitute the first token of each text segment with [BOS] token to evaluate the performance of IntactKV.
MMLU.
We evaluate MMLU following the original MMLU implementation [^6] https://github.com/hendrycks/test/pull/13 for 0-shot and 5-shot tasks. We note that when using Vicuna, it is considered more appropriate to fit the input sequences into the Vicuna system prompt. However, the original MMLU implementation does not use the Vicuna system prompt for Vicuna models. In our experiments on Vicuna models, we find that naively fitting the original MMLU prompt into the Vicuna system prompt will harm the final accuracy. Since prompt engineering is out of scope for this paper, we choose to follow the original evaluation setting that does not use the Vicuna system prompt for MMLU evaluation on Vicuna models.
Common Sense Reasoning Tasks.
For the seven zero-shot common sense reasoning tasks, we adopt the open-sourced lm-evaluation-harness [^7] https://github.com/EleutherAI/lm-evaluation-harness library for evaluation. Similar to PPL evaluation, to assess the performance of IntactKV, we prepend [BOS] token to the beginning of each input sequence. For the evaluation of Vicuna models, we also follow the evaluation protocol in lm-evaluation-harness and do not use a system prompt.
MT-bench.
MT-bench employs a GPT-4 model to score the generated content. In our experiments, we find that the score given by GPT-4 can vary for the same content even when the generation temperature of GPT-4 is set to 0. Besides, content generation for the writing and roleplay category has a relatively high generation temperature of 0.7, which also results in variations in the final score. To faithfully assess the performance of the quantized model, we run the content generation process of each model 3 times with random seeds 42, 43, and 44. We report the mean score of three trials as the final score in Table 5 and Table 6. Also, we note that GPT-4-Turbo has been shown to be smarter than GPT-4888https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard, and in our experiments, we find that GPT-4-Turbo can give more stable score than GPT-4. Therefore, we evaluate the generation results on MT-bench with the latest gpt-4-0125-preview API (i.e., GPT-4-Turbo) provided by OpenAI to further reduce variations in the final score.
Appendix F More Experiment Results
F.1 PPL Results on WikiText2
As shown in Table 7, AWQ+IntactKV consistently improves AWQ for every model on the WikiText2 dataset and outperforms OmniQuant for five out of seven models. We note that OmniQuant uses the WikiText2 dataset as a calibration set, while AWQ uses Pile dataset, which may give OmniQuant a bonus for PPL evaluation on the WikiText2 dataset.
F.2 MMLU Results
We provide INT3-g128 quantization results on MMLU in Table 8, and INT4-g128 quantization results on MMLU in Table 9. For INT3-g128 quantization, AWQ+IntactKV consistently improves AWQ in every experiment setting and outperforms OmniQuant for nine out of ten settings. For INT4-g128 quantization, AWQ+IntactKV leads to relatively less improvement over AWQ on the MMLU benchmark compared with INT3-g128 quantization, but still outperforms AWQ in nine out of ten experiment settings, and performs on par with OmniQuant on MMLU.
F.3 Commonsense QA Results
We conduct experiments on seven zero-shot commonsense QA tasks for the Vicuna family with both INT3-g128 and INT4-g128 quantization. The results are shown in Table 10. For INT3-g128 quantization, AWQ+IntactKV significantly surpasses all baselines in all experiment settings. For INT4-g128 quantization, AWQ+IntactKV consistently improves AWQ and outperforms OmniQuant in four out of five experiment settings, demonstrating the necessity for keeping an intact KV cache for pivot tokens.
Appendix G Experiment Details of KV Cache Quantization
We apply asymmetric per-head dynamic quantization to the KV cache. When combined with IntactKV, we still keep IntactKV in FP16 while the KV cache of other tokens in low bit, which only induces negligible memory overhead compared with quantizing the entire KV cache to low bits since IntactKV only contains the KV cache of a few tokens.
Appendix H Links to Officially Released LLMs
We provide download links to some officially released LLMs used in our experiments in Table 11.
Figure 6: Magnitude of the output activations and attention map in LLaMA-7B.
Figure 7: Magnitude of the output activations and attention map in LLaMA-13B.
Figure 8: Magnitude of the output activations and attention map in LLaMA-30B.
Figure 9: Magnitude of the output activations and attention map in LLaMA-65B.
Figure 10: Magnitude of the output activations and attention map in LLaMA-2-7B.
Figure 11: Magnitude of the output activations and attention map in LLaMA-2-13B.
Figure 12: Magnitude of the output activations and attention map in LLaMA-2-70B.
Figure 13: Magnitude of the output activations and attention map in Vicuna-v1.3-7B. The tokens before the red dashed line correspond to the Vicuna system prompt.
Figure 14: Magnitude of the output activations and attention map in Vicuna-v1.3-13B. The tokens before the red dashed line correspond to the Vicuna system prompt.
Figure 15: Magnitude of the output activations and attention map in Vicuna-v1.3-33B. The tokens before the red dashed line correspond to the Vicuna system prompt.
Figure 16: Magnitude of the output activations and attention map in Vicuna-v1.5-7B. The tokens before the red dashed line correspond to the Vicuna system prompt.
Figure 17: Magnitude of the output activations and attention map in Vicuna-v1.5-13B. The tokens before the red dashed line correspond to the Vicuna system prompt.